helppp plzzProblem 1-25 Traditional and Contribution Format Income Statements (LO1-6] Milden Company is a merchandiser that plans to sel
Answer #1

Answer:

1)The contribution format income statement for the next quarter is the following way:

Milden company
  Contribution margin income statement
For the next quarter
Particulars Amount Amount
Sales

= total sale of units*selling price per unit

=39,000*65 = 25,35,000

$25,35,000
Less variable Expenses:
variable cost of goods sold(39,000*32) $12,48,000
Variable selling expenses(39,000*4) $1,56,000
Variable selling commission(25,35,000*5/100) $1,26,750
Total variable Expenses $15,30,750
Contribution margin(sales-total variable costs)(25,35,000-15,30,750) $10,04,250
Less Fixed Expenses :
Advertising Expenses $1,88,000
Shipping expense $42,000
Administrative expense $98,000
Insurance expenses $10,800
Depreciation expense $68,000
Total Fixed Expenses $4,06,800
Net income (contribution margin -total fixed expenses)(10,04,250-4,06,800) $5,97,450

2)The Traditional format income statement for the next quarter is the following way:

   Milden company
   Traditional Format income statement
For the next quarter
Particulars Amount Amount
Sales

= total sale of units*selling price per unit

=39,000*65 = 25,35,000

$25,35,000
Cost of goods sold(39,000*32) $12,48,000
Gross margin $12,87,000
Selling and administrative expenses:
Advertising Expenses $1,88,000
Shipping expense(42,000+selling expenses[39,000*4]) $1,98,000
Administrative expense $98,000
Insurance expenses $10,800
Depreciation expense $68,000
sales commission(25,35,000*5/100) $1,26,750
Total selling and administrative expenses $6,89,550
Net operating income(gross margin-Total selling and administrative expenses)

=12,87,000-6,89,550

=5,97,450

$5,97,450